Identify and describe the five most important ratios for a potential shareholder to assess prior to making a stock purchase in a company. Explain the reasoning for your selection. Illustrate your discussion using numerical examples.
Answer #1

The five most important ratios for a potential shareholder to assess prior to making a stock purchase in a company are:

  1. Price to earnings ratio (P/E) – This is the ratio of the company's current market price to the company's earnings per share. This ratio has been selected by me as it is important to see how much investors are willing to pay for the single dollar of earnings of the company. The higher ratio means that investors are willing to pay more while lower the ratio investors are willing to pay less for the earnings of the company. For instance let us assume that current price of Apple’s share is $10 per share and its earnings per share is $2. Thus its P/E ratio = 10/2 = 5. This can be compared to P/E figures of other similar stocks like Google, Microsoft etc. to determine if the stock of Apple is commanding a premium or not.
  2. Price to book value (P/B) – Book value per share is computed by dividing the assets of the company with the number of shares. Price to book value is calculated by dividing the current market price with the book value per share of the company. The P/B ratio indicates the amount that will be left for shareholders in the event the company goes bankrupt. For instance assume that the price per share is $10 and book value per share is $1 Thus P/B ratio = 10/12 = 0.83. This is less than 1 and indicates that the stock is undervalued as value of assets on the company's books is more than the value the market is assigning to the company.
  3. Debt to equity ratio – This ratio indicates leverage and is computed by dividing total liabilities of the company by the equity shareholders capital of the company. The lower the ratio the better is its financial risk situation. For example assume that for Apple its debt is $100 and its equity is $200 and so its debt equity ratio = 100/200 = 0.5. This should be compared to debt equity ratio of peer companies of Apple.
  4. Return on equity – This ratio measure how much return the company is giving to its shareholders and shows the ability of the management to generate profits. Return on equity = net income/equity shareholder capital. Assume that for Apple the net income is $100 and shareholder capital is $2000. Thus return on equity = 100/2000 = 5%. This figure should be compared to Apple’s historical return on equity as well as to return on equity numbers of its peers.
  5. Dividend yield – This ratio is computed by dividing the annual dividend per share of the company with the share price of the company. This ratio helps in determining how much the company pays out dividend to its shareholders per year. Assume that for Apple’s stock the annual dividend per share is $10 and its share price is $100. Thus dividend yield is 10/100 = 10%. During uncertain times and during volatile times investors would usually park their investments in shares of those companies that have high dividend yield as it will compensate them for the absence of capital gains.